OpenStudy (anonymous):

tan ( pi/2- theta) tan theta=1

OpenStudy (jdoe0001):

ok.. .thanks for sharing that

OpenStudy (anonymous):

thank you they're asking to verify the identity.

OpenStudy (jdoe0001):

notice -> http://www.gradeamathhelp.com/image-files/trigonometric-identities-pythagorean.gif <--- the 1st pythagorean identity there

OpenStudy (jdoe0001):

ahemm rather the 2nd one listed there

OpenStudy (jdoe0001):

ohh smokes... .wait a sec... I got the wrong set

OpenStudy (jdoe0001):

http://1.bp.blogspot.com/-57vrQkzCnVA/TxfkOuTBs6I/AAAAAAAABWU/R91jwBewUyE/s1600/cofunction+identities.png <--- there rather, we'll use the 3rd one listed there, from the trig cofunctions

OpenStudy (anonymous):

is this how the answer should be Cot theta tan theta=1 cos theta / sin theta times sin theta / cos theta=1 1=1

OpenStudy (jdoe0001):

well... .kinda.. one sec

OpenStudy (anonymous):

sure

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ tan\left(\frac{\pi}{2}-\theta\right)}}=cot(\theta)\qquad {\color{olive}{ cot(\theta)}}=\cfrac{cos(\theta)}{sin(\theta)} \\ \quad \\ \quad \\ tan\left(\frac{\pi}{2}-\theta\right)tan(\theta)=1\implies {\color{brown}{ cot(\theta)}}tan(\theta)=1 \\ \quad \\ {\color{olive}{ \cfrac{\cancel{ cos(\theta) }}{\cancel{ sin(\theta) }}}} \cdot \cfrac{\cancel{ sin(\theta) }}{\cancel{ cos(\theta) }}=1\)

OpenStudy (anonymous):

wht is tht ?

OpenStudy (jdoe0001):

hmmm anything you find confusing? there's an assumption you know trig btw

OpenStudy (anonymous):

if you are a beginner,then \[\tan \left( \frac{ \pi }{ 2 } - \theta \right)=\frac{ \sin \left( \frac{ \pi }{ 2}-\theta \right) }{ \cos \left( \frac{ \pi }{ 2 } -\theta \right) }\] \[=\frac{ \sin \frac{ \pi }{ 2 }\cos \theta-\cos \frac{ \pi }{ 2} \sin \theta }{ \cos \frac{ \pi }{ 2 }\cos \theta+\sin \frac{ \pi }{ 2 } \sin \theta }\] \[(\sin \frac{ \pi }{ 2 }=1,\cos \frac{ \pi }{ 2 }=0)\] \[=\frac{ \cos \theta }{ \sin \theta }\]